1.     Drive the AD (Aggregate Demand) curve using the following: IS curve is given as Y = 20XX-100i, LM1 is Y= 1000+25i (when P = 1) and LMis Y = 500+25i (when P =2), where XX is the last two digits of your student ID number. Show the derivation in (interest rate-income) and (price level-income) spaces. (You may insert a snapshot of the graphs if drawn manually).



1
Expert's answer
2021-05-07T09:53:57-0400

IS curve:

 "Y = 20XX -100i"

"XX=30"

"Y =2030 -100i"

The equilibrium output is determined where IS = LM.

At price 1,

"2030 \u2212100i=1000 +\u200925i"

"125i=1030"

"i=8.24"

"Y=2030-100(8.24)"

"Y=2030-824"

"Y=1206"


At price 2,

"2030\u2212100i=500 + 25i"

"125i=1530"

"I=12.24"

"Y=2030 \u2212100(12.24)"

"Y=2030 \u22121224"

"Y=806"



Interest rate income:

Y1i"=1206"

Y2i"=806"

Deviation"=1206-806"

"=400"

Price level income:

Y1p"=1000+25(1)"

"=1025"

Y2p"=500+25(2)"

"=550"

Deviation"=1025-550"

"=475"